1.) Based on the N/Z ratio, size of the nuclide, and knowledge of stable nuclides, what mode of decay is most likely for the following nuclides?

a. Potassium-40

b. Chlorine-41

c. Oxygen-20

d. Thorium-224

e. Carbon-15

2) Oxygen-16 is one of the most stable nuclides. The mass of an oxygen-16 atom is 15.994915. Calculate the mass defect (or the binding energy) of oxygen forming from neutrons and protons in kJ/mol.
Answer #1

1. C. Beta minus. (Beta-)

D. Alpha into Rn-220

A. All 3 types of beta decay

B. Beta-

E. Beta -

2. Mass defect = sum of the mass of the nucleons - mass of the nucleus

O has 8 protons = 1.00728×8 = 8.05824

8 neutrons= 1.00867×8= 8.06936

Total mass of nucleons = 16.1276

Mass defect = 16.1276 - 15.994915 = 0.132685